pat is 6 years older than his friend Gary. Four years ago Gary's age was 7 more than half of Pat's age. What are their present age?
Answer by stanbon(75887)

pat is 6 years older than his friend Gary. Four years ago Gary's age was 7 more than half of Pat's age. What are their present age?
---
Equations:
p = g+6
(g-4)= (1/2)p + 7
---
Substitute for "p" and solve for "g":
g-4 = (1/2)(g+6)+7
2g-8 = g+13
g = 21 (Gary's age now)
---
Solve for "p":
p = 21+6
p = 27 (Pat's age now)
================================
Cheers,
Stan H.